OpenStudy (anonymous):

inverse laplace of F(s) = ln|(s-2)/(s+2)|

OpenStudy (abb0t):

Take the integral as you did in the previous problem you asked for.

OpenStudy (anonymous):

So I tried going backwards from the last problem like exactly step by step and it gave me an answer but entirely wrong haha

OpenStudy (abb0t):

\[\iota ^-1\int\limits_{s}^{∞} F(s)ds \]

OpenStudy (anonymous):

but i dont have f(t) i have F(s)

OpenStudy (anonymous):

crap haha

hartnn (hartnn):

do u know this property ? \(\huge L^{-1}[F(s)]=\frac{-1}{t}L^{-1}[\frac{d}{ds}F(s)]\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we know F(s)

hartnn (hartnn):

so what did u get d/ds (F(s)) ?

OpenStudy (anonymous):

but how do we take LT inverse of it

hartnn (hartnn):

what did u get d/ds (F(s)) ?

hartnn (hartnn):

first use. ln A/B = ln A-ln B

OpenStudy (anonymous):

4/(s^2 - 4)

OpenStudy (anonymous):

okay i think i got it! :)

hartnn (hartnn):

u get something of this form, A/(s+2) + B/(s-2)

hartnn (hartnn):

or u can directly use hyperbolic function, i think...

OpenStudy (anonymous):

wait no i dont have it

OpenStudy (anonymous):

so i got 4/(s^2 - 4) which is 4/1 * 1/(s^2 - (2)^2)

hartnn (hartnn):

is the answer in terms of hyperbolic function or exponential function ?

hartnn (hartnn):

because that will decide next step...

hartnn (hartnn):

ln|(s-2)/(s+2)| = ln (s-2) - ln(s+2) derivative will be = 1/s-2 - 1/ s+2 now the inverse LT of this is easy.

OpenStudy (abb0t):

pellet son! i 4got laplace transform.

OpenStudy (anonymous):

oh okay

hartnn (hartnn):

so, did u get the answer ?

OpenStudy (anonymous):

no :(

hartnn (hartnn):

what u got ? and whats the actual answer ?

OpenStudy (anonymous):

answer: -(2sinh(2t))/t

OpenStudy (anonymous):

oh so the 2 is from the 4

hartnn (hartnn):

yeah, u know the Laplace of sinh t , right ?

hartnn (hartnn):

-1/t came from the property 4/(s^2-2^2) = 2 *[2/(s^2-2^2)] ---> 2sinh t got it ?

OpenStudy (anonymous):

yeah sure sounds good